OpenStudy (anonymous):

which is half of 2+2????

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Not...

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

rsrs

sammixboo (sammixboo):

Is this a real question, or a question to throw people off?

OpenStudy (ipwnbunnies):

To throw you off.

OpenStudy (anonymous):

a matter that fell in the vestibular here in Brazil, logical reasoning

OpenStudy (ipwnbunnies):

We don't have the mental capacity to understand the depth of the answer.

OpenStudy (anonymous):

por curiosidade por isso postei, mas vamos lá, a metada de 2+2 é..... metade de 2 é = à 1 +2 é = à 3 Resposta é 3.....

OpenStudy (anonymous):

out of curiosity why I posted, but come on, to half of 2 +2 is ..... half of 2 is 1 +2 is = to the = 3 Response is 3 .....

OpenStudy (ipwnbunnies):

Your question was ambiguous. Lack of parentheses.

OpenStudy (anonymous):

But really this is logical reasoning, so dropped the vestibular .... had to interpret the question and solve